40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 2 12:51:52 2011, kiwamu, Update, LSC, PRMI locking : plan 
    Reply  Mon May 2 13:30:13 2011, kiwamu, Update, SUS, PRMI locking prep: PRM oplev 
    Reply  Mon May 2 18:16:24 2011, kiwamu, Update, LSC, PRMI locking : status update 
    Reply  Tue May 3 03:53:20 2011, kiwamu, Update, LSC, PRMI locking : done PRMI_locking.png
       Reply  Tue May 3 16:47:55 2011, kiwamu, Update, LSC, PRMI locking : details APtable.png
Message ID: 4596     Entry time: Mon May 2 13:30:13 2011     In reply to: 4595
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: PRMI locking prep: PRM oplev 

I had a quick look at PRM optical lever.

The He-Ne beam is still successfully coming out from the chamber and I could guide it to the QPD by using steering mirrors.

But the beam size looks too big for the QPD. We should slide the lens which is standing before the injection to get a moderately smaller beam size at the QPD.

Quote from #4595

 - activation of PRM oplev

ELOG V3.1.3-